I been working on these problems PLEASE HELP!

1. The body mass index, I, can be used to determine an individual’s risk for heart disease. An index less than 25 indicates a low risk. The body mass index is given by the formula, or model. I=700W/H2, where W=weight, in pounds, and H=height, in inches. Francis weights 160 pounds and stands 74 inches tall. What is his approximate body mass index? Find an inequality describing all weight W that Francis can have and be in the low-risk category.
Francis’ approximate body mass index is___
(Round to the nearest whole number
Find and inequality describing all weights W that Francis can have and be in the low-risk category,
{W l W< (approximately)__lb}
Round down to the nearest whole pound)


2. Bayside Insurance offers two health plans. Under plan A, Giselle would have to pay the first $120 of her medical bills, plus 25% of the rest. Under plan B, Giselle would pay the first $220, but only 20% of the rest. For what amount of medical bills will plan B save Giselle money? Assume she has over $220 in bills/
Giselle would save with plan B if she had more than $__in bills.
(Round to the nearest whole number)


3. In order to achieve maximum results from aerobic excise. One should maintain one’s heart rate at a certain level. A 40 year-old woman with a resting heart rate of 60 beats per minute should keep her heart rate between 132 and 156 beats per minute while exercising. She checks her pulse for 10 second while exercising. What should the number of beats be?
The number of beats in 10 sec. should be between ___and ___beats.
(Type the smaller number in the first answer box and the large number in the second answer box.

4. ¾(2x-2)>15
Use set-builder notation to describe the complete solution.
{x l x__ __}
Simplify your answer

Answer by stanbon(48558)

1. The body mass index, I, can be used to determine an individual’s risk for heart disease. An index less than 25 indicates a low risk. The body mass index is given by the formula, or model. I=700W/H^2, where W=weight, in pounds, and H=height, in inches. Francis weights 160 pounds and stands 74 inches tall. What is his approximate body mass index?
---
I = 700*160/74^2 = 20.4529
------------------------------------
Find an inequality describing all weight W that Francis can have and be in the low-risk category.
700W/H^2 < 25
700W/74^2 < 25
W < 74^2*25/700
W < 195.57 lbs.
----------------------

2. Bayside Insurance offers two health plans. Under plan A, Giselle would have to pay the first $120 of her medical bills, plus 25% of the rest. Under plan B, Giselle would pay the first $220, but only 20% of the rest. For what amount of medical bills will plan B save Giselle money? Assume she has over $220 in bills/
Giselle would save with plan B if she had more than $__in bills.
(Round to the nearest whole number)
---
Plan A: cost = 120 + 0.25(x-120) where "x" is the total medical bill cost.
Plan B: cost = 220 + 0.20(x-220)
---
Inequality
Plan B < Plan A ??
220 + 0.20(x-220) < 120 + 0.25(x-120)
220 + 0.20x - 44 < 120 + 0.25x - 30
100 -14 < 0.05x
0.05x > 86
x > $1720 (if the bill is over $1720 Plan B will save her money)
-------------------------------------------


3. In order to achieve maximum results from aerobic excise. One should maintain one’s heart rate at a certain level. A 40 year-old woman with a resting heart rate of 60 beats per minute should keep her heart rate between 132 and 156 beats per minute while exercising. She checks her pulse for 10 second while exercising. What should the number of beats be?
The number of beats in 10 sec. should be between ___and ___beats.
(Type the smaller number in the first answer box and the large number in the second answer box.
Per minute: 132 < beats per minute < 156
Per 10 seconds : 132/6 < beats per 10 seconds < 156/6
22 < beats per 10 seconds < 26
-------------------------------------
4. ¾(2x-2)>15
Use set-builder notation to describe the complete solution.
(3/4)(2x-2) > 15
2x-2 > (4/3)*15 = 20
2x > 22
x > 11
{x l x > 11}
=======================================
Cheers,
Stan H.